a man had a bag of sweets he gave one to his son and 1/7 of the remaining and what he was left he gave to his daughter two sweet and 1/7 of the remaining sweets.the two children found that they have same number of sweets how many sweets are there in original bag
Answer by mathstutor494(120)

Let there be "x" sweets in original bag
Gave 1 to son so remaining are x-1 sweets
1/7 of remaining sweets = (x-1)*(1/7)=> (x-1)/7

Hence total sweets given to son
1+((x-1)/7)
(x+6)/7 ...............................(1)
Sweets remaining in the bag after giving to son
x-((x+6)/7)
(6x-6)/7
Out of this he gave 2 sweets to daughter so remaining sweets are
((6x-6)/7)-2
=(6x-20)/7
Further, he gave 1/7 from remaining lot to his daughter...
((6x-20)/7)*(1/7)
=(6x-20)/49
So total sweets with daughter
((6x-20)/49)+2
=(6x+78)/49 ..............................(2)
Since the problem specifies that the two children found that they have same number of sweets, so expressions (1) and (2) are equal.
Accordingly
(x+6)/7 = (6x+78)/49
Solving for x gives value of 36
Ans : 36 sweets there in original bag